OpenStudy (pegasii):

x · 1 +x/1= _____. A.) 1 B.) x C.) 1 + x D.) 2x

OpenStudy (fifciol):

\[x\cdot 1+ \frac{x}{1} = ?\] What happens to some arbitrary number, when you multiply, or divide it by one?

OpenStudy (pegasii):

stays as 1 ?

sam (.sam.):

@Pegasii Nope, the term stays the same, \(a \times 1 = a\), \(\frac{a}{1}=a\)

sam (.sam.):

So.. any progress?

OpenStudy (pegasii):

it equals answer b then?

sam (.sam.):

How can it be B? You have \[x \times 1+ \frac{x}{1} \\ \\ \\ \\\] \[x+x=?\]